X company issues 5,000 ten-year 8% debentures of rs 100 each at per redeemable at 5% at premium. The cost of issue is 2%. What is the cost of debt capital?
1
Expert's answer
2021-09-06T07:59:05-0400

Solution:

The cost of debt capital (Kd) = "I +( \frac{RV - NP}{n}) \div ( \frac{RV + NP}{2})"


Where: I = Interest on Debenture = 8% x 100 = 8

           NP = Net Proceeds = (105% x 100) – (2% of 105% x 100) = 105 – 2.1 = 102.9

           RV = Redemption Value = 100

             n = Period of Debenture = 10 years


Kd = "8 +( \frac{100 - 102.9}{10}) \div ( \frac{100 + 102.9}{2})"


Kd = "\frac{8 - 0.29}{101.45} = \frac{7.71}{101.45} = 0.076 \; or\; 7.6\%"


The cost of debt capital (Kd) = "7.6\%"